# On additive error approximations to ##\##BQP

### v1

@@ 3.1 Additive Error Approximations to #BQP
If we take $l>\frac{w_{r}(|x|)}{1-c}+1$ then the error is less than $1/2$ and by rounding to the nearest integer, a number in $r(x,\cdot)$ can always be returned. This takes $l=O\left(\frac{w_{r}(|x|)}{1-c}\right)$ extra qubits and thus $l$, and the size of $V^{r^{\prime}}_{x}$, will be upper bounded by a polynomial in $|x|$ whenever $1-c=\Omega(\text{poly}(|x|)^{-1})$.
Additive approximations to #BQP are BQP-hard
Prior to proving Theorem 2.3, we introduce a different version of #BQP that allows for the $c^{r}$ and $s^{r}$ functions to depend on the input $x$, rather than just its size, as utilizing this version will greatly simplify the proof.
Definition 9 (#BQP${}_{(c^{r},s^{r})}$).
A relation $r\subset J\times\mathbb{Z}^{+}$ is in #BQP if and only if there exists a uniformly generated family of quantum circuit $V^{r,n}$ and a polynomial $q^{r}$ such that
$r(x)=[N_{\geq c^{r}(x)}(\mathcal{V}_{x}^{r,|x|})..N_{\geq s^{r}(x)}(\mathcal{V}_{x}^{r,|x|})]\,,$
where $c^{r},s^{r}:\{0,1\}^{*}\to[0,1]$ are efficiently computable functions satisfying $c^{r}(x)-s^{r}(x)>1/q^{r}(|x|)$.
We show in Appendix A that Definition 9 is equivalent to Definition 5 in order to ensure that this new definition is not more powerful. This yields the following lemma.
Lemma 3.1.
#BQP${}_{(c^{r},s^{r})}$ = #BQP
We are now in a position to prove Theorem 2.3.
Proof of Theorem 2.3.
The goal here is to show that if we can solve a #BQP${}_{(c^{r},s^{r})}$ relation $r$ on a series of polynomially many inputs $x_{i}$ to additive error, with normalizations $u_{i}=2^{w(|x_{i}|)}$, then we can solve an input $x$ of an instance of the Average Accept Probability of Verifier Circuits problem in Definition 3 with parameters $c(|x|)$ and $s(|x|)$. In particular, let $x\mapsto x_{i}=y_{i}x$ where $y_{i}$ is a bit string which encodes the correct completeness and soundness parameters to use for the corresponding input. Choosing a fixed length prefix-free encoding for $y_{i}$ ensures uniqueness of each concatenated bit string, and each $x_{i}$ has the same number of bits, which will be necessary to avoid repeated inputs. Each input $x_{i}$ to the #BQP${}_{(c^{r},s^{r})}$ relation we are constructing will have a verifier circuit $V^{|x_{i}|}$ whose corresponding acceptance operator will be $\mathcal{V}_{x}$, the same operator as that of the Average Accept Probability of Verifier Circuits input. The $y_{i}$ part of the bit string will then determine the $c_{i}$ and $s_{i}$ interval.
Then we begin by partitioning the interval $[0,1]$ into $M=\text{poly}(|x|)$ subintervals $[s_{i},c_{i}]$, each of size $1/\text{poly}(|x|)$, corresponding to the different completeness and soundness parameters specified by $y_{i}$. Specifically, we consider the subintervals $[s_{i},c_{i}]$ where $c_{i}=(M-i)/M$ and $s_{i}=c_{i}-1/4M$ for $11\leq i\leq M-$ 1. For $c_{M}$ and $s_{M}$ the oracle will always return $2^{w(|x|)}$. Then querying the #BQP${}_{(c^{r},s^{r})}$ oracle for each problem input $x_{i}$ returns an estimate
$\hat{n}_{i}=N_{\geq c_{i}}+\delta_{i}+\varepsilon_{i}\,,$ (16)
with $|\delta_{i}|\in[0,N_{[s_{i},c_{i}]}]$ and $|\varepsilon_{i}|\in[0,\epsilon 2^{w(|x_{i}|)}]$. Letting $\hat{n}_{0}=0$, we define the two quantities $\hat{\chi}_{i}=\hat{n}_{i}-\hat{n}_{i-1}$ and $\Delta_{i}=c_{i}+1/2M$. Then we have
$=\sum_{i=1}^{M}\Delta_{i}N_{[c_{i},c_{i-1}]}+\frac{1}{M}(\delta_{i}+\varepsilon_{i})\,.$ (18)
We will now show that this quantity approximates the trace of the associated acceptance operator $\mathcal{V}_{x}$ of the BQP problem. First, note that
$|\Delta_{i}N_{[c_{i},c_{i-1}]}-\text{Tr}_{[c_{i},c_{i-1}]}(\mathcal{V}_{x^{\prime}})|\leq\frac{1}{2M}N_{[c_{i},c_{i-1}]}\,,$ (19)
which then implies that
$\bigg{|}\sum_{i=1}^{M}\Delta_{i}N_{[c_{i},c_{i-1}]}-\text{Tr}(\mathcal{V}_{x})\bigg{|}\leq\frac{2^{w(|x|)}}{2M}\,,$ (20)
where we have used the fact that $2^{w(|x_{i}|)}=2^{w(|x|)}$ for all $i$, or rather, the size of the witness subspace for each input $x_{i}$ is the same size as that of $x$ even though the classical inputs may vary in size.
$\frac{1}{M}\sum_{i=1}^{M}\delta_{i}\leq\frac{2^{w(|x|)}}{M},\quad\text{and}\quad\sum_{i=1}^{M}\varepsilon_{i}\leq\epsilon 2^{w(|x|)}\,,$ (21)
where the first inequality holds from the fact that $[s_{i},c_{i}]\cap[s_{j},c_{j}]=\emptyset$ for all $i\neq j$. Setting $\epsilon=1/M$, we add the errors from each contribution and find that
$\bigg{|}\sum_{i=1}^{M}\Delta_{i}\hat{\chi}_{i}-\text{Tr}(\mathcal{V})\bigg{|}\leq\frac{2^{w(|x|)}}{M}+\frac{2^{w(|x|)}}{M}+\frac{2^{w(|x|)}}{2M}=\frac{5}{2}\frac{2^{w(|x|)}}{M}\,.$ (22)
Choosing $M>5/4(c(|x|)-s(|x|))$ suffices to distinguish ebetween the accepting and rejecting cases of Definition 3, a BQP-complete problem. Thus, returning additive approximations to #BQP problems is BQP-hard.
Classical additive approximations to #BQP relations
Here we prove Theorem 2.4.
Proof of Theorem 2.4.
@@
Given a #BQP relation $r$ and an input $x$ with a corresponding verifier circuit $V_{x}^{r}$, set $n=w_{r}(|x|)$ as the number of witness qubits, $a=a_{r}(|x|)$ as the number of ancilla qubits, $T=t_{r}(|x|)$ as the size of the circuit in the $\{H,S,\text{Toffoli}\}$ gate set, and $h=h_{r}(|x|)$ as the number of Hadamard gates. Let $V_{x}^{r}=g_{T}\cdots g_{1}$ in the chosen gate set and let $\mathcal{V}$ be the corresponding verifier operator. First note that because of the completeness and soundness parameters we have that
$N(\mathcal{V})_{\geq c}-(1-c)2^{n}\leq\text{Tr}[\mathcal{V}]\leq N(\mathcal{V})_{\geq s}+s2^{n}\,.$ (23)
@@
$=$ $\sum_{x,y,z_{j}}\langle 0^{a}x|g_{1}^{\dagger}|z_{1}\rangle\!\langle z_{1}|g_{2}^{\dagger}|z_{2}\rangle\cdots\langle z_{T-1}|g_{T}^{\dagger}|1y\rangle\!\langle 1y|g_{T}|z_{T}\rangle\!\langle z_{T}|g_{T-1}|z_{T+1}\rangle\cdots\langle z_{2(T-1)}|g_{1}|0^{a}x\rangle$ (25)
$=$ $\frac{1}{2^{h}}\sum_{x,y,z_{j}}\langle 0^{a}x|q_{1}^{\dagger}|z_{1}\rangle\!\langle z_{1}|q_{2}^{\dagger}|z_{2}\rangle\cdots\langle z_{T-1}|q_{T}^{\dagger}|1y\rangle\!\langle 1y|q_{T}|z_{T}\rangle\!\langle z_{T}|q_{T-1}|z_{T+1}\rangle\cdots\langle z_{2(T-1)}|q_{1}|0^{a}x\rangle.$ (26)
@@
$\langle 0^{a}x|q_{1}^{\dagger}|z_{1}\rangle\!\langle z_{1}|q_{2}^{\dagger}|z_{2}\rangle\cdots\langle z_{T-1}|q_{T}^{\dagger}|1y\rangle\!\langle 1y|q_{T}|z_{T}\rangle\!\langle z_{T}|q_{T-1}|z_{T+1}\rangle\cdots\langle z_{2(T-1)}|q_{1}|0^{a}x\rangle$ (27)
@@
$\frac{r}{2^{h}}=\text{Tr}[\mathcal{V}]\,.$ (28)
Next, it follows from Ref. [1] that we can additively approximate $r$ with $u=2^{N^{*}}$ where $N^{*}=|p|$ is the number of bits needed to describe the paths $p$. Since $|x|=n$, $|y|=n+a-1$, and $|z_{j}|=n+a$ for all $j$, it follows that the number of bits to describe the path is $N^{*}=2T(n+a)-(a+1)$. Now, it follows by approximating $r$ and dividing by $2^{h}$, we get an additive approximation $\tilde{\mathcal{V}}$ to $\text{Tr}[\mathcal{V}]$ with $u=2^{N^{*}}/2^{h}$ we have
$N_{\geq c}(\mathcal{V})-(1-c)2^{n}-\epsilon 2^{N^{*}-h}\leq\text{Tr}[\tilde{\mathcal{V}}]\leq N_{\geq s}(\mathcal{V})+s2^{n}+\epsilon 2^{N^{*}-h}$ (29)
@@
$N_{\geq c}(\mathcal{V})-\epsilon^{\prime}2^{N^{*}-h}\leq\text{Tr}[\tilde{\mathcal{V}}]\leq N_{\geq s}(\mathcal{V})+\epsilon^{\prime\prime}2^{N^{*}-h}$ (30)
@@
In this section we prove Theorems 2.5 and 2.6, noting that the proof of Theorem 2.5 follows closely to that of Theorem 2.3.
Additive approximations to #BQP with logarithmic ancillas is DQC1-complete.
Proof of Theorem 2.5.
Again, just as in the proof of Theorem 2.3, the goal here is to show that if we can solve a #BQP${}_{(c^{r},s^{r})}$ relation $r$ on a polynomial number of inputs $x_{i}$ to additive error, with normalizations $u_{i}=2^{w(|x_{i}|)}$, then we can solve the Average Accept Probability of Verifier Circuits problem in Definition 3. However, there are a few notable differences here. First, the notion of uniform generation of quantum circuits changes when considering DQC1 so rather than starting from Definition 5, we use the alternative definition of #BQP in Definition 8. Its equivalence to #BQP${}_{(c^{r},s^{r})}$ follows exactly that of the proof in Appendix A. Additionally, here we consider only a sublcass of #BQP problems, namely those whose verifier circuits $D^{|x|}$ require at most a logarithmic number of ancilla qubits. Provided each step in our reduction to the Average Accept Probability of Verifier Circuits problem, which is DQC1-complete in this regime, does not use more than $O(\log(|x|)$ ancilla qubits, then we will show that additively approximating any relation $r$ in this subclass of #BQP is DQC1-hard.
The techniques used in the proof Theorem 2.3 will be the exact same here, so rather than restating the proof, we will just ensure each step preserves the size of the ancilla space. In fact, the only part of the proof that can change the ancilla space is in the reduction of Definition 8 to Definition 9 in Appendix A, in which we employ the quantum singular value transformation. However, this step only requires a single ancilla qubit to implement. Thus, provided that the verifier circuits $D^{|x|}$ require only $O(\log(|x|)$ ancilla qubits, then a polynomial number of calls to the #BQP${}_{(c^{r},s^{r})}$ oracle to additive error allows us to additively approximate any DQC1 problem. ∎
Proof of Theorem 2.6.
Let $V_{x}^{r}$ be a given verifier circuit for a #BQP relation $r$ on an input $x$ with $w_{r}(|x|)$ qubits and $a_{r}(|x|)=O(\log(w_{r}(|x|)))$ ancilla qubits with completeness and soundness parameters $c(|x|)$ and $s(|x|)$ and an associated verifier operator $\mathcal{V}_{x}$. Then one can always construct a block encoding $A\coloneqq\bra{1}V_{x}^{r}(\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})$, and then make use of Ref. [13, Theorem 19] to implement a singular value threshold projector on $A$. This allows us to construct a new block encoding $A_{\Phi}=\bra{1}V_{\Phi}(\ket{0}\ket{+}\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})$, for some $\Phi\in\mathbb{R}^{p}$, and where $V_{\Phi}$ is given by the circuit in Fig. 1. The new encoding $A_{\Phi}$ now has singular values $\tilde{\varsigma}_{i}$ which are $\varepsilon$-close to 00 or $1$ depending on whether the singular values $\varsigma_{i}$ of $A$ are below or above a threshold $t\in(0,1)$.
To construct the new block encoding $A_{\Phi}$, define a function on the interval $[-1,1]$ which takes the value 00 on the interval $[-t,t]$ and $1$ elsewhere. A polynomial approximation $P$ can be constructed such that
@@
Additionally, it can be shown that this polynomial approximation is even and has degree $p=O(\log(1/\varepsilon)/\Delta)$ using techniques similar to that of Ref. [13, Corollary 16]. Finally, according to Ref. [13, Corollary 11], given the polynomial approximation $P$, one can construct a vector $\Phi\in\mathbb{R}^{p}$ which applies $P$ to the singular values of $A$, resulting in the new encoding $A_{\Phi}=P(A)$. The implementation of $P(A)$ is based on quantum signal processing techniques, and requires only one ancilla qubit with $p$ applications of $V_{x}^{r}$ and $(V_{x}^{r})^{\dagger}$ and $O(p)$ other elementary gates, as shown in Fig. 1.
Thus, given the verifier circuit $V_{x}^{r}$, we define the new verifier circuit $V_{\Phi}$ in Fig. 1 that implements the singular value threshold projector protocol on its block encoding $A$, setting $t=(s(|x|)+c(|x|))/2$ and $\Delta=(c(|x|)-s(|x|))/2$, to obtain the new encoding $A_{\Phi}=\bra{1}V_{\Phi}(\ket{0}\ket{+}\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})$ with associated acceptance operator
$\mathcal{V}_{\Phi}$ $=A_{\Phi}^{\dagger}A_{\Phi}\,,$ (32)
[…]
$=\sum_{i}\tilde{\varsigma}_{i}\ket{\psi_{i}}\bra{\psi_{i}}\,,$ (34)
@@
$N_{\geq c(|x|)}(\mathcal{V}_{x})-(2\varepsilon-\varepsilon^{2})2^{w_{r}(|x|)}\leq\text{Tr}[\mathcal{V}_{\Phi}]\leq N_{\geq s(|x|)}(\mathcal{V}_{x})+\varepsilon^{2}2^{w_{r}(|x|)}\,.$ (35)
@@
Finally, the trace of this amplified acceptance operator can be additively estimated to normalization $u(|x|)=2^{w_{r}(|x|)}$ and precision $\epsilon$ via the same algorithm as in the proof of Theorem 2.1 using $O(1/\epsilon^{2})$ repetitions of $V_{\Phi}$. This outputs an estimate $\tilde{X}$ satisfying
$N_{\geq c(|x|)}(\mathcal{V}_{x})-(2\varepsilon-\varepsilon^{2}+\epsilon)2^{w_{r}(|x|)}\leq\tilde{X}\leq N_{\geq s(|x|)}(\mathcal{V}_{x})+(\varepsilon^{2}+\epsilon)2^{w_{r}(|x|)}\,,$ (36)
@@
with probability greater than $3/4$. Thus by taking $\varepsilon=O(\log(\text{poly}(\epsilon^{\prime})^{-1}))$ such that both $2\varepsilon-\varepsilon^{2}$ and $\varepsilon^{2}$ are less than or equal to $\epsilon$, an additive error estimate to normalization $2^{w_{r}(|x|)}$ and precision $\epsilon^{\prime}=2\epsilon$ can be obtained. Then, noting that $\Delta=(c(|x|)-s(|x|))/2=\Omega(\text{poly}(|x|)^{-1})$, an additive error estimate to normalization $u(|x|)=2^{w(|x|)}$ and to precision $\epsilon^{\prime}$ can be obtained using $O(\text{poly}\left(1/\epsilon^{\prime}\right))$ runs of $O(\log(\text{poly}(1/\epsilon^{\prime}))\cdot\text{poly}(|x|))$ size quantum circuits. Then repeating this procedure $\log(1/\delta^{\prime})$ times and taking the median of estimates can improve the success probability to $1-\delta^{\prime}$. Importantly since the circuits used in this algorithm require only a logarithmic overhead of $a_{r}(|x|)+2=O(\log(w_{r}(|x|)))$ pure ancilla qubits, they can be computed using DQC1 circuits.
Figure 1: The circuit $V_{\Phi}$ implementing the quantum singular value transformation $A_{\Phi}=P(A)$ given in Ref. [13] using a degree-$p$, even polynomial approximation. The block encodings $V_{x}^{r}$ and $V_{x}^{r\dagger}$ are applied alternately followed by applications of the reflection operators. In particular, controlled on the witness register being in the image of $\Pi$, the ancilla qubit is flipped, then a single qubit Pauli $Z$ rotation is applied to angles $\phi_{i}$, where the particular angles are classically determined according to the function being approximated by the polynomial. The final control gate performs an $X$ operation on the top qubit if and only if all of its control qubits are in the 00 state. The top $a_{r}(|x|)+2$ qubits are ancillas and the bottom $w_{r}(|x|)$ qubits are the witness qubits.

### v2

@@ 4.1 Additive-Error Approximations to #BQP
If we take $l>\frac{w_{r}(|x|)}{1-c}+1$ then the error is less than $1/2$ and by rounding to the nearest integer, a number in $r(x,\cdot)$ can always be returned. This takes $l=O\left(\frac{w_{r}(|x|)}{1-c}\right)$ extra qubits and thus $l$, and the size of $V^{{}^{\prime}r,(|x|)}$, will be upper bounded by a polynomial in $|x|$ whenever $1-c=\Omega(\text{poly}(|x|)^{-1})$.
We additionally show that efficient classical algorithms for returning additive-error approximations to #BQP functions exist, though they can yield larger normalization factors.
See 3.1
Proof.
@@
Given a #BQP relation $r$ and an input $x$ with a corresponding verifier circuit $V^{r,(|x|)}$, set $n=w_{r}(|x|)$ as the number of witness qubits, $a=a_{r}(|x|)$ as the number of ancilla qubits, $T=t_{r}(|x|)$ as the size of the circuit in the $\{H,S,\text{Toffoli}\}$ gate set, and $h=h_{r}(|x|)$ as the number of Hadamard gates. Let $V^{r,(|x|)}=g_{T}\cdots g_{1}$ in the chosen gate set and let $\mathcal{V}_{x}^{r}$ be the corresponding verifier operator. First note that because of the completeness and soundness parameters we have that
$N_{\geq c}(\mathcal{V}_{x}^{r})-(1-c)2^{n}\leq\text{Tr}[\mathcal{V}_{x}^{r}]\leq N_{\geq s}(\mathcal{V}_{x}^{r})+s2^{n}\,.$ (20)
@@
$=\sum_{y,z_{j}}\langle 0^{a},x|g_{1}^{\dagger}|z_{1}\rangle\!\langle z_{1}|g_{2}^{\dagger}|z_{2}\rangle\cdots\langle z_{T-1}|g_{T}^{\dagger}|1,y\rangle\!\langle 1,y|g_{T}|z_{T}\rangle\!\langle z_{T}|g_{T-1}|z_{T+1}\rangle\cdots\langle z_{2(T-1)}|g_{1}|0^{a},x\rangle$
$=\frac{1}{2^{h}}\sum_{y,z_{j}}\langle 0^{a},x|q_{1}^{\dagger}|z_{1}\rangle\!\langle z_{1}|q_{2}^{\dagger}|z_{2}\rangle\cdots\langle z_{T-1}|q_{T}^{\dagger}|1,y\rangle\!\langle 1,y|q_{T}|z_{T}\rangle\!\langle z_{T}|q_{T-1}|z_{T+1}\rangle\cdots\langle z_{2(T-1)}|q_{1}|0^{a},x\rangle.$
@@
$\langle 0^{a},x|q_{1}^{\dagger}|z_{1}\rangle\!\langle z_{1}|q_{2}^{\dagger}|z_{2}\rangle\cdots\langle z_{T-1}|q_{T}^{\dagger}|1,y\rangle\!\langle 1,y|q_{T}|z_{T}\rangle\!\langle z_{T}|q_{T-1}|z_{T+1}\rangle\cdots\langle z_{2(T-1)}|q_{1}|0^{a},x\rangle$ (22)
@@
$\frac{r}{2^{h}}=\text{Tr}[\mathcal{V}_{x}^{r}]\,.$ (23)
Next, it follows from Ref. [1] that we can additively approximate $r$ with $u=2^{N^{*}}$ where $N^{*}=|p|$ is the number of bits needed to describe the paths $p$. Since $|x|=n$, $|y|=n+a-1$, and $|z_{j}|=n+a$ for all $j$, it follows that the number of bits to describe the path is $N^{*}=2T(n+a)-(a+1)$. Now, it follows by approximating $r$ and dividing by $2^{h}$, we get an additive approximation $\tilde{\mathcal{V}}_{x}^{r}$ to $\text{Tr}[\mathcal{V}_{x}^{r}]$ with $u=2^{N^{*}}/2^{h}$ we have
$N_{\geq c}(\mathcal{V}_{x}^{r})-(1-c)2^{n}-\epsilon 2^{N^{*}-h}\leq\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]\leq N_{\geq s}(\mathcal{V}_{x}^{r})+s2^{n}+\epsilon 2^{N^{*}-h}$ (24)
@@
$N_{\geq c}(\mathcal{V}_{x}^{r})-\epsilon^{\prime}2^{N^{*}-h}\leq\text{Tr}[\tilde{\mathcal{V}}_{x}^{r}]\leq N_{\geq s}(\mathcal{V}_{x}^{r})+\epsilon^{\prime\prime}2^{N^{*}-h}$ (25)
@@
Here we prove that additive-error approximations to a subclass of #BQP functions, namely those whose verifier circuits require only a logarthimic number of ancilla qubits, is DQC1-complete with normalization scaling exponential in the dimension of the witness space.
See 3.2
Proof.
In order to show that BPPA $\subseteq$ DQC1, we prove there exists a polynomial-time quantum circuit, using only a logarithmic number of ancilla qubits, which simulates $A$. In particular, let $D^{r,(x)}$ be a given verifier circuit for a #BQP relation $r$ on an input $x$ with $w_{r}(|x|)$ qubits and $a_{r}(|x|)=O(\log(w_{r}(|x|)))$ ancilla qubits with completeness and soundness parameters $c(|x|)$ and $s(|x|)$ and an associated verifier operator $\mathcal{D}_{x}^{r}$. Then one can always construct a block encoding $U\coloneqq\bra{1}D^{r,(x)}(\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})$, and then make use of Ref. [13, Theorem 19] to implement a singular value threshold projector on $U$. This allows us to construct a new block encoding $U_{\Phi}=\bra{1}D_{\Phi}(\ket{0}\ket{+}\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})$, for some $\Phi\in\mathbb{R}^{p}$, and where $D_{\Phi}$ is given by the circuit in Fig. 1. The new encoding $U_{\Phi}$ now has singular values $\tilde{\varsigma}_{i}$ which are $\varepsilon$-close to 00 or $1$ depending on whether the singular values $\varsigma_{i}$ of $U$ are below or above a threshold $t\in(0,1)$.
Figure 1: The circuit $D_{\Phi}$ implementing the quantum singular value transformation $U_{\Phi}=P(U)$ given in Ref. [13] using a degree-$p$, even polynomial approximation. The block encodings $D^{r,(x)}$ and $(D^{r,(x)})^{\dagger}$ are applied alternately followed by applications of the reflection operators. In particular, controlled on the witness register being in the image of $\Pi$, the ancilla qubit is flipped, then a single qubit Pauli-$Z$ rotation is applied to angles $\phi_{i}$, where the particular angles are classically determined according to the function being approximated by the polynomial. The final control gate performs an $X$ operation on the top qubit if and only if all of its control qubits are in the 00 state. The top $a_{r}(|x|)+2$ qubits are ancillas and the bottom $w_{r}(|x|)$ qubits are the witness qubits.
To construct the new block encoding $U_{\Phi}$, define a function on the interval $[-1,1]$ which takes the value 00 on the interval $[-t,t]$ and $1$ elsewhere. A polynomial approximation $P$ can be constructed such that
@@
Additionally, it can be shown that this polynomial approximation is even and has degree $p=O(\log(1/\varepsilon)/\Delta)$ using techniques similar to that of Ref. [13, Corollary 16]. Finally, according to Ref. [13, Corollary 11], given the polynomial approximation $P$, one can construct a vector $\Phi\in\mathbb{R}^{p}$ which applies $P$ to the singular values of $U$, resulting in the new encoding $U_{\Phi}=P(U)$. The implementation of $P(U)$ is based on quantum signal processing techniques, and requires only one ancilla qubit with $p$ applications of $D^{r,(x)}$ and $(D^{r,(x)})^{\dagger}$ and $O(p)$ other elementary gates, as shown in Fig. 1.
Thus, given the verifier circuit $D^{r,(x)}$, we define the new verifier circuit $D_{\Phi}$ in Fig. 1 that implements the singular value threshold projector protocol on its block encoding $U$, setting $t=(s(|x|)+c(|x|))/2$ and $\Delta=(c(|x|)-s(|x|))/2$, to obtain the new encoding $U_{\Phi}=\bra{1}D_{\Phi}(\ket{0}\ket{+}\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})$ with associated acceptance operator
$\mathcal{D}_{\Phi}$ $=U_{\Phi}^{\dagger}U_{\Phi}\,,$ (27)
$=(\bra{0}\bra{+}\bra{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})D_{\Phi}^{\dagger}|1\rangle\langle 1|D_{\Phi}(\ket{0}\ket{+}\ket{0}^{\otimes a_{r}(|x|)}\otimes\mathbb{I}_{W})\,,$
$=\sum_{i}\tilde{\varsigma}_{i}\ket{\psi_{i}}\bra{\psi_{i}}\,,$
@@
$N_{\geq c(|x|)}(\mathcal{D}_{x}^{r})-(2\varepsilon-\varepsilon^{2})2^{w_{r}(|x|)}\leq\text{Tr}[\mathcal{D}_{\Phi}]\leq N_{\geq s(|x|)}(\mathcal{D}_{x}^{r})+\varepsilon^{2}2^{w_{r}(|x|)}\,.$ (28)
@@
Finally, the trace of this amplified acceptance operator can be additively estimated to normalization $u(|x|)=2^{w_{r}(|x|)}$ and precision $\epsilon$ via the same algorithm as in the proof of Theorem 3.1 using $O(1/\epsilon^{2})$ repetitions of $D_{\Phi}$. This outputs an estimate $\tilde{X}$ satisfying
$N_{\geq c(|x|)}(\mathcal{D}_{x}^{r})-(2\varepsilon-\varepsilon^{2}+\epsilon)2^{w_{r}(|x|)}\leq\tilde{X}\leq N_{\geq s(|x|)}(\mathcal{D}_{x}^{r})+(\varepsilon^{2}+\epsilon)2^{w_{r}(|x|)}\,,$ (29)
@@
with probability greater than $3/4$. Thus by taking $\varepsilon=O(\log(\text{poly}(\epsilon^{\prime})^{-1}))$ such that both $2\varepsilon-\varepsilon^{2}$ and $\varepsilon^{2}$ are less than or equal to $\epsilon$, an additive error estimate to normalization $2^{w_{r}(|x|)}$ and precision $\epsilon^{\prime}=2\epsilon$ can be obtained. Then, noting that $\Delta=(c(|x|)-s(|x|))/2=\Omega(\text{poly}(|x|)^{-1})$, an additive error estimate to normalization $u(|x|)=2^{w(|x|)}$ and to precision $\epsilon^{\prime}$ can be obtained using $O(\text{poly}\left(1/\epsilon^{\prime}\right))$ runs of $O(\log(\text{poly}(1/\epsilon^{\prime}))\cdot\text{poly}(|x|))$ size quantum circuits. Then repeating this procedure $\log(1/\delta^{\prime})$ times and taking the median of estimates can improve the success probability to $1-\delta^{\prime}$. Importantly since the circuits used in this algorithm require only a logarithmic overhead of $a_{r}(|x|)+2=O(\log(w_{r}(|x|)))$ pure ancilla qubits, they can be computed using DQC1 circuits. Therefore BPPA $\subseteq$ DQC1.
On the other hand, to show that DQC1 $\subseteq$ BPPA, we use similar techniques as those in the proof of Theorem 3.1. We show that querying $A$ for a #BQP${}_{(c^{r},s^{r})}$ relation $r$ with $a_{r}(|x|)=O(\log(w_{r}(|x|)))$ ancilla qubits on a polynomial number of inputs $x_{i}$ to additive error, with normalizations $u_{i}=2^{w(|x_{i}|)}$, allows us to solve any DQC1 decision problem as given in Definition 4. In fact, the Average Accept Probability for Verifier Circuits problem in Definition 7 reduces to the definition of DQC1 in Definition 4 when only a logarithmic number of ancilla qubits are allowed. A subtle difference here is that we should like the circuits used to be generated on an input-by-input basis, as is done in the definition of DQC1. Rather than starting from Definition 6 as before, we modify it to allow for input-dependent verifier-circuit generation, with the acceptance operator $\mathcal{D}_{x}^{r}$ corresponding to the circuit $D^{r,(x)}$ for relation $r$ and input $x$.
Establishing this fact, the remainder of the proof is identical to the proof of Theorem 3.1, replacing $\mathcal{V}_{x}^{r}$ with $\mathcal{D}_{x}^{r}$. First, partition the interval $[0,1]$ into $M=\mathcal{O}(\text{poly}(|x|))$ subintervals $[s_{i},c_{i}]$, each of size $\Omega(\text{poly}(|x|)^{-1})$, corresponding to the different completeness and soundness parameters specified by the fixed-length prefix-free encoding $y_{i}$. Specifically, we consider the subintervals $[s_{i},c_{i}]$ where $c_{i}=(M-i)/M$ and $s_{i}=c_{i}-1/4M$ for $11\leq i\leq M-$ 1. Then querying $A$ for each problem input $x_{i}$ returns an estimate
$\hat{n}_{i}=N_{\geq c_{i}}+\delta_{i}+\varepsilon_{i}\,,$ (30)
with $\delta_{i}\in[0,N_{[s_{i},c_{i}]}]$ and $|\varepsilon_{i}|\in[0,\epsilon 2^{w(|x_{i}|)}]$. Letting $\hat{n}_{0}=0$, we define the two quantities $\hat{\chi}_{i}=\hat{n}_{i}-\hat{n}_{i-1}$ and $\Delta_{i}=c_{i}+1/2M$. Then we have
$=\sum_{i=1}^{M}\Delta_{i}N_{[c_{i},c_{i-1}]}+\sum_{i=1}^{M-1}(\delta_{i}+\varepsilon_{i})+\underbrace{\frac{1}{2M}(\delta_{M}+\epsilon_{M})}_{=0}\,.$
where the last term vanishes as a result of the fact that for the interval $[s_{M},c_{M}]$, we need not query the oracle and instead always return $2^{w(|x|)}$, meaning $\delta_{M},\epsilon_{M}=0$.
This quantity approximates the trace of the associated acceptance operator $\mathcal{D}_{x}^{r}$ of DQC1 in Definition 4. In particular, note that
$|\Delta_{i}N_{[c_{i},c_{i-1}]}-\text{Tr}_{[c_{i},c_{i-1}]}(\mathcal{D}_{x}^{r})|\leq\frac{1}{2M}N_{[c_{i},c_{i-1}]}\,,$ (32)
which then implies that
$\bigg{|}\sum_{i=1}^{M}\Delta_{i}N_{[c_{i},c_{i-1}]}-\text{Tr}(\mathcal{D}_{x}^{r})\bigg{|}\leq\frac{2^{w(|x|)}}{2M}\,.$ (33)
$\frac{1}{M}\sum_{i=1}^{M}\delta_{i}\leq\frac{2^{w(|x|)}}{M},\quad\text{and}\quad\frac{1}{M}\sum_{i=1}^{M}\varepsilon_{i}\leq\epsilon 2^{w(|x|)}\,,$ (34)
where the first inequality holds from the fact that $[s_{i},c_{i}]\cap[s_{j},c_{j}]=\emptyset$ for all $i\neq j$. In the second inequality we have used the fact that $2^{w(|x_{i}|)}=2^{w(|x|)}$ for all $i$, or rather, the size of the witness subspace for each input $x_{i}$ is the same size as that of $x$ even though the classical inputs may vary in size. Setting $\epsilon=1/M$, we add the errors from each contribution and find that
$\bigg{|}\sum_{i=1}^{M}\Delta_{i}\hat{\chi}_{i}-\text{Tr}(\mathcal{D}_{x}^{r})\bigg{|}\leq\frac{2^{w(|x|)}}{M}+\frac{2^{w(|x|)}}{M}+\frac{2^{w(|x|)}}{2M}=\frac{5}{2}\frac{2^{w(|x|)}}{M}\,.$ (35)
Choosing $M>5/(c(|x|)-s(|x|))$ suffices to distinguish between the accepting and rejecting cases of Definition 4. Therefore, DQC1 $\subseteq$ BPPA. ∎